# Numerical investigation of the interior geometry of semiclassical evaporating spherical charged black holes

### v1

@@ VI.3 $g_{1}$ metric
$\widetilde{v}(v)\equiv\stackrel{{\scriptstyle[}}{{0}}]{v}{\int}\kappa_{+}^{v}(v)dv,\hfill\widetilde{u}(u)\equiv\stackrel{{\scriptstyle[}}{{0}}]{u}{\int}\kappa_{+}^{u}(u)du$ (VI.25)
where $\kappa_{+}^{u}(u)$ was defined in Eq. (IV.13). These coordinates satisfy the following relations:
$\frac{d\widetilde{v}}{dv}=\kappa_{+}^{v}(v),\hfill\frac{d\widetilde{u}}{du}=\kappa_{+}^{u}(u)$ (VI.26)
@@
$g_{uv}$ $=g_{\widetilde{u}\widetilde{v}}\frac{\partial\widetilde{u}}{\partial u}\frac{\partial\widetilde{v}}{\partial v}$ (VI.27)
$=\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)\frac{1}{2\kappa_{+}^{v}(v)^{2}}\cdot\left(\kappa_{+}^{v}(v)\kappa_{+}^{u}(u)\right)$ (VI.28)
@@
$ds^{2}=\frac{\kappa_{+}^{u}(u)}{\kappa_{+}^{v}(v)}fdu\;dv+r^{2}d\Omega^{2}$ (VI.29)
@@
$S=\log\left(-rg_{uv}\right)=\log\left(-\frac{1}{2}rf\right)+\log\left(\kappa_{+}^{u}(u)\right)-\log\left(\kappa_{+}^{v}(v)\right)$ (VI.31)
To get the function $R=r^{2}$, we recall the definition of $r_{*}$ given in Eq. (VI.19)
$r_{*}=\frac{\widetilde{u}+\widetilde{v}}{2\kappa_{+}^{v}(v)}$ (VI.32)
@@
$\frac{\widetilde{u}(u)+\widetilde{v}(v)}{2\kappa_{+}^{v}(v)}=r_{*}=r-M+\frac{1}{2\kappa_{+}}\log\left|\frac{r-r_{+}}{M-r_{+}}\right|-\frac{1}{2\kappa_{-}}\log\left|\frac{r-r_{-}}{M-r_{-}}\right|$ (VI.33)
The $u$ derivative of $r$ is then (using Eq. (VI.20)):
$r_{,u}=r_{,\widetilde{u}}\widetilde{u}_{,u}=f\frac{\kappa_{+}^{u}(u)}{2\kappa_{+}^{v}(v)}$ (VI.34)
@@
$R_{,u}=2rr_{,u}=rf\frac{\kappa_{+}^{u}(u)}{\kappa_{+}^{v}(v)}$ (VI.35)
@@
$S=\log\left(-rg_{uv}\right)=\log\left(-\frac{1}{2}rf\right)+\log\left(\kappa_{+}^{u}(u)\right)-\log\left(\kappa_{+}^{v}(v)\right)$ (VI.36)
therefore,
$S_{,u}=\left(\frac{\partial}{\partial r}\log\left(-rf\right)\right)r_{,u}+\frac{1}{\kappa_{+}^{u}(u)}\frac{d}{du}\kappa_{+}^{u}(u)$ (VI.37)
@@
$\frac{\partial}{\partial r}\log\left(-rf\right)=\frac{1-\frac{Q^{2}}{r^{2}}}{rf}$ (VI.38)
@@
$\frac{d}{du}\left(\kappa_{+}^{u}(u)\right)=\frac{\partial}{\partial M}\kappa_{+}^{u}(u)\frac{d}{du}M^{u}(u)+\frac{\partial}{\partial Q}\kappa_{+}^{u}(u)\frac{d}{du}Q^{u}(u)$ (VI.39)
Finally, we show that if u+v=0,u+v=0,then $\widetilde{u}+\widetilde{v}=0$ and therefore $r=M(v)$. Choosing a specific value $v=v_{0}$, from $u+v=0$ we obtain that $u=-v_{0}.$ We then get from Eq. (VI.25) that:
$\widetilde{v}+\widetilde{u}=\stackrel{{\scriptstyle[}}{{0}}]{v_{0}}{\int}\kappa_{+}^{v}(v)dv+\stackrel{{\scriptstyle[}}{{0}}]{-v_{0}}{\int}\kappa_{+}^{u}(u)du$ (VI.40)
@@
$\widetilde{v}+\widetilde{u}$ $=\stackrel{{\scriptstyle[}}{{0}}]{v_{0}}{\int}\kappa_{+}^{v}(v)dv+\stackrel{{\scriptstyle[}}{{0}}]{-v_{0}}{\int}\kappa_{+}^{v}(-u)du$ (VI.42)
$=\stackrel{{\scriptstyle[}}{{0}}]{v_{0}}{\int}\kappa_{+}^{v}(v)dv-\stackrel{{\scriptstyle[}}{{0}}]{v_{0}}{\int}\kappa_{+}^{v}(w)dw=0$ (VI.43)
(In the last equality we changed the integration variable w=uw=-u.) We therefore conclude that when $u+v=0$, $\widetilde{u}+\widetilde{v}=0$ and therefore (as seen in subsection VI.2) $r=M(v)$.
VI.4 Initial hypersurface and the range of the $v$ and $u$ coordinates
@@
At the point $v=0$ on the initial hypersurface, namely the point marked A in Figure VI.1, $\frac{u+v}{2}=\frac{u}{2}=T_{0}$ and therefore $u=2T_{0}$ there. At the point $v=\Delta v$ on the initial hypersurface, namely the point marked B, $\frac{u+v}{2}=\frac{u+\Delta v}{2}=T_{0}$ and therefore $u=2T_{0}-\Delta v$. Therefore, the choice of the value of $T_{0}$ determines the range of $u$: $u$ runs from $2T_{0}-\Delta v$ to $2T_{0}$. To summarize, the point A is located at $v=0,u=2T_{0}$, the point B at $v=\Delta v,u=2T_{0}-\Delta v$; therefore the point C is at $v=\Delta v,u=2T_{0}-\Delta v$.
The domain of the numerical integration is the grey equilateral right triangle shown in Figure VI.1. The hypotenuse of the triangle is the initial hypersurface, and the two sides are defined by the two null rays $u=2T_{0}$ and $v=\Delta v$.
Figure VI.1: v and u coordinates
@@
We now present the results of running the numerical code. Recall that we set our units such that $M_{0}=L_{0}=1$. As a default, we set $du=dv=\frac{1}{100}$, meaning there are hundred grid points per 1 unit of $u$ or $v$. We set the initial-moment parameter $T_{0}=-30$ as a default (unless stated otherwise). Recall that the range of $v$ is usually from $v=0$ to $v=\Delta v=900$, and correspondingly the range of $u$ is from $u=2T_{0}-\Delta v=-960$ to $u=2T_{0}=-60$.
@@
First we will present the results of running with no source terms (which mean that $T_{uv}=T_{\theta\theta}=0$) in section VII. We will present and explore the resulting quantities $R,S$, $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$. In region 2 we will compare the evolving metric versus that of the corresponding local RN metric. Then we turn to explore region 3, namely the near-IH region. We will graphically display both the numerical result and the corresponding analytical approximation from section IV. Then we shall also display the deviation from the analytical approximation caused primarily by finite $\varepsilon_{0}$.
@@
Then we will present the result of running with all sources set to 1 in section VIII. We shall again display the quantities of the analytical approximation and their corresponding numerical results, as well as the deviations from the analytical approximation.
@@
Then in the next stage in section IX we will analyze the consequences of running with finite $T_{0}$ (the value of $T=\frac{u+v}{2}$ on the initial surface) and finite $du$ – by running with several different values of these parameters and comparing the results.
@@
Afterwards in section X we will display that the detail of sources doesn’t matter, only the values of $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$. We will do it by running with 3 different combinations of the parameters $Z_{R0}$$Z_{Rr}$$Z_{S0}$$Z_{Sr}$ chosen such that they all lead to the same $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$. We display the metric functions $R$ and $S$ in these configurations, showing that they are the same in those 3 configurations.
@@
Lastly we display the results of running with three other different combinations of source parameters and show the validity of the analytical approximation in all those runs in section XI
VII Numerical results - run with no sources
@@
We start by presenting the results of running with zero source terms, $Z_{R0}=$$Z_{Rr}$$Z_{S0}=$$Z_{Sr}=$0=0, as well as $\delta Q=0$, for $\frac{Q}{M}=0.8$. In this case, $T_{uu}=0$ throughout the whole spacetime, as already mentioned in subsection II.8 above. We will first look at the entire domain of integration (including all three regions). Then we shall have a more focused look: first on region 2, and then on region 3. In region 2 we will compare the results with the local RN metric. In region 3 we will compare the metric functions to the analytical approximation from section IV. We will then analyze the deviation from the analytical approximation.
VII.1 Overview
Figure VII.1: R
@@ VII.1 Overview
As you can observe in Figure VII.1a, showing $R(v)$, $R$ starts at a value of $R_{+}\equiv r_{+}^{2}=2.56$ in region 1, drifts down because of the negative $M_{,v}$ (still in region 1); then it “falls down” during the active region (region 2) – as it does in RN – and ends near $R_{-}\equiv r_{-}^{2}=0.16$ near IH, in the beginning of region 3. Then it drifts slightly downwards with a small negative $R_{,v}$ according to Eq. (IV.40).
The second panel VII.1b, displaying $R(u)$, shows a slightly different behavior. Because $T_{uu}=0$ (unlike $T_{vv}$ which is not zero), $R$ stays almost constant in region 1, $R\approx r_{+}^{2}\left(M(v=900),Q(v=900)\right)=0.551535$, and also in region 3 where $R\approx r_{-}^{2}\left(M(v=900),Q(v=900)\right)=0.034471$. It only varies significantly in region 2, similar to the RN case. In this special case where $J_{u}=0$ and $T_{uu}=0$ the geometry is the charged Vaidya solution, therefore along an incoming null ray with $v=const$ all quantities behave exactly like in the corresponding RN.
Figure VII.2: S
@@
The following figure VII.2 shows the behavior of $S$. We first focus on $S(v)$ in panel VII.2a. In region 1, $S_{,v}=\kappa_{+}(v)$, and hence $S$ increases towards region 2 (but this is hard to notice in the figure because of the vertical scale, which mostly shows region 3. The behavior in regions 1 and 2 is better seen in Figure VII.8a below). In region 2, $S$ peaks and then goes down as $S_{,v}=-\kappa_{-}(v)$ on approaching the near-IH region. In region 3, $S$ keeps drifting down over the long stretch from $v=100$ until $v=900$ for the same reason. $S$ eventually reaches very negative values of -4000 and even below. Therefore, $e^{S}\approx 0$ to a very good approximation.
Panel VII.2b is very similar to panel VII.2a, and shows $S$ as a function of $u$, along the other ray $v=900$.
Figure VII.3: $\widetilde{T}_{uu}$ on last v,u rays
@@
Next, we show in Figure VII.3 that $\widetilde{T}_{uu}=0$ in the entire domain of integration. The small deviation from zero (primarily seen in region 2) is due to numerical truncation error.
Figure VII.4: $\widetilde{T}_{vv}$ on last u ray
@@
We continue to Figure VII.4 which shows both $\widetilde{T}_{vv}$ and the quantity
$2M_{,v}-2\frac{1}{r}QQ_{,v}$ (VII.1)
which is the expected value of $\widetilde{T}_{vv}$ according to Eq. (II.109). The two curves overlap, demonstrating that $\widetilde{T}_{vv}$ matches its expected value. In region 2, $\widetilde{T}_{vv}(v)$ goes up from -0.0004 to +0.0004 (due to the rapid change in $r$ in the above expression $2\frac{1}{r}QQ_{,v}$). It then continues to drift in region 3 (though more slowly) as all factors $M_{,v}$ , $r$,$Q$, $Q_{,v}$ appearing in VII.1 slowly drift there with increasing $v$.
Figure VII.5: $\widetilde{T}_{vv}(u)$ on last v ray
Next we see in Figure VII.5 that $\widetilde{T}_{vv}(u)$ is constant in regions 1 and 3, and it only varies in region 2. This is because $M_{,v}$ and $Q_{,v}$ and $Q$ depend only on $v$, hence these quantities are constant along the last $v$ ray. Therefore the only change in the expression VII.1 comes from changing $r$. But $r$ doesn’t change in regions 1 and 3, it only changes in region 2.
Figure VII.6: $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ on initial hypersurface
@@
(b) $\widetilde{T}_{vv}-\left(2M_{,v}-2\frac{1}{r}QQ_{,v}\right)$ on initial hypersurface
@@
In Figure VII.6 we elaborate on the fluxes $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ on the initial hypersurface (the red line in Figure VI.1), namely $T=T_{0}$. Panel VII.6a compares $\widetilde{T}_{vv}$ with its theoretically expected value $2M_{,v}-2\frac{1}{r}QQ_{,v}$. We display the small difference between these two quantities in Figure VII.6b, which shows the error reaching at most $2\cdot 10^{-5}$. We see that the computed value matches the expected value well. Figure VII.6c, shows that up to numerical errors ($2\cdot 10^{-10}$, mainly roundoff), $T_{uu}=0$ on the initial hypersurface line, as it should.
VII.2 Region 2: comparison with the local RN approximation

### v2

@@ 6.3 $g_{1}$ metric
$\widetilde{v}(v)\equiv\mathrel{{\mathop{\int}\limits_{0}^{v}}}\kappa_{+}^{v}(v)dv,\hfill\widetilde{u}(u)\equiv\mathrel{{\mathop{\int}\limits_{0}^{u}}}\kappa_{+}^{u}(u)du$ (6.25)
where $\kappa_{+}^{u}(u)$ was defined in Eq. (4.13). These coordinates satisfy the following relations:
$\frac{d\widetilde{v}}{dv}=\kappa_{+}^{v}(v),\hfill\frac{d\widetilde{u}}{du}=\kappa_{+}^{u}(u)$ (6.26)
@@
$g_{uv}$ $=g_{\widetilde{u}\widetilde{v}}\frac{\partial\widetilde{u}}{\partial u}\frac{\partial\widetilde{v}}{\partial v}$ (6.27)
$=\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)\frac{1}{2\kappa_{+}^{v}(v)^{2}}\cdot\left(\kappa_{+}^{v}(v)\kappa_{+}^{u}(u)\right)$ (6.28)
@@
$ds^{2}=\frac{\kappa_{+}^{u}(u)}{\kappa_{+}^{v}(v)}fdu\;dv+r^{2}d\Omega^{2}$ (6.29)
@@
$S=\log\left(-rg_{uv}\right)=\log\left(-\frac{1}{2}rf\right)+\log\left(\kappa_{+}^{u}(u)\right)-\log\left(\kappa_{+}^{v}(v)\right)$ (6.31)
To get the function $R=r^{2}$, we recall the definition of $r_{*}$ given in Eq. (6.19)
$r_{*}=\frac{\widetilde{u}+\widetilde{v}}{2\kappa_{+}^{v}(v)}$ (6.32)
@@
$\frac{\widetilde{u}(u)+\widetilde{v}(v)}{2\kappa_{+}^{v}(v)}=r_{*}=r-M+\frac{1}{2\kappa_{+}}\log\left|\frac{r-r_{+}}{M-r_{+}}\right|-\frac{1}{2\kappa_{-}}\log\left|\frac{r-r_{-}}{M-r_{-}}\right|$ (6.33)
The $u$ derivative of $r$ is then (using Eq. (6.20)):
$r_{,u}=r_{,\widetilde{u}}\widetilde{u}_{,u}=f\frac{\kappa_{+}^{u}(u)}{2\kappa_{+}^{v}(v)}$ (6.34)
@@
$R_{,u}=2rr_{,u}=rf\frac{\kappa_{+}^{u}(u)}{\kappa_{+}^{v}(v)}$ (6.35)
@@
$S_{,u}=\left(\frac{\partial}{\partial r}\log\left(-rf\right)\right)r_{,u}+\frac{1}{\kappa_{+}^{u}(u)}\frac{d}{du}\kappa_{+}^{u}(u)$ (6.36)
@@
$\frac{\partial}{\partial r}\log\left(-rf\right)=\frac{1-\frac{Q^{2}}{r^{2}}}{rf}$ (6.37)
@@
$\frac{d}{du}\left(\kappa_{+}^{u}(u)\right)=\frac{\partial}{\partial M}\kappa_{+}^{u}(u)\frac{d}{du}M^{u}(u)+\frac{\partial}{\partial Q}\kappa_{+}^{u}(u)\frac{d}{du}Q^{u}(u)$ (6.38)
Finally, we show that if u+v=0,u+v=0,then $\widetilde{u}+\widetilde{v}=0$ and therefore $r=M(v)$. Choosing a specific value $v=v_{0}$, from $u+v=0$ we obtain that $u=-v_{0}.$ We then get from Eq. (6.25) that:
$\widetilde{v}+\widetilde{u}=\mathrel{{\mathop{\int}\limits_{0}^{v_{0}}}}\kappa_{+}^{v}(v)dv+\mathrel{{\mathop{\int}\limits_{0}^{-v_{0}}}}\kappa_{+}^{u}(u)du$ (6.39)
@@
$\widetilde{v}+\widetilde{u}$ $=\mathrel{{\mathop{\int}\limits_{0}^{v_{0}}}}\kappa_{+}^{v}(v)dv+\mathrel{{\mathop{\int}\limits_{0}^{-v_{0}}}}\kappa_{+}^{v}(-u)du$ (6.41)
$=\mathrel{{\mathop{\int}\limits_{0}^{v_{0}}}}\kappa_{+}^{v}(v)dv-\mathrel{{\mathop{\int}\limits_{0}^{v_{0}}}}\kappa_{+}^{v}(w)dw=0$ (6.42)
(In the last equality we changed the integration variable w=uw=-u.) We therefore conclude that when $u+v=0$, $\widetilde{u}+\widetilde{v}=0$ and therefore (as seen in subsection 6.2) $r=M(v)$.
6.4 Initial hypersurface and the range of the $v$ and $u$ coordinates
@@
At the point $v=0$ on the initial hypersurface, namely the point marked A in Figure 6.1, $\frac{u+v}{2}=\frac{u}{2}=T_{0}$ and therefore $u=2T_{0}$ there. At the point $v=\Delta v$ on the initial hypersurface, namely the point marked B, $\frac{u+v}{2}=\frac{u+\Delta v}{2}=T_{0}$ and therefore $u=2T_{0}-\Delta v$. Therefore, the choice of the value of $T_{0}$ determines the range of $u$: $u$ runs from $2T_{0}-\Delta v$ to $2T_{0}$. To summarize, the point A is located at $v=0,u=2T_{0}$, the point B at $v=\Delta v,u=2T_{0}-\Delta v$; therefore the point C is at $v=\Delta v,u=2T_{0}-\Delta v$.
The domain of the numerical integration is the grey equilateral right triangle shown in Figure 6.1. The hypotenuse of the triangle is the initial hypersurface, and the two sides are defined by the two null rays $u=2T_{0}$ and $v=\Delta v$.
Figure 6.1: v and u coordinates
@@
We now present the results of running the numerical code. Recall that we set our units such that $M_{0}=L_{0}=1$. As a default, we set $du=dv=\frac{1}{100}$, meaning there are hundred grid points per 1 unit of $u$ or $v$. We set the initial-moment parameter $T_{0}=-30$ as a default (unless stated otherwise). Recall that the range of $v$ is usually from $v=0$ to $v=\Delta v=900$, and correspondingly the range of $u$ is from $u=2T_{0}-\Delta v=-960$ to $u=2T_{0}=-60$. Recall that we set $\varepsilon_{0}=\frac{1}{1000}.$
@@
First we will present the results of running with no source terms (which mean that $T_{uv}=T_{\theta\theta}=0$) in section 7. We will present and explore the resulting quantities $R,S$, $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$. In region 2 we will compare the evolving metric versus that of the corresponding local RN metric. Then we turn to explore region 3, namely the near-IH region. We will graphically display both the numerical result and the corresponding analytical approximation from section 4. Then we shall also display the deviation from the analytical approximation caused primarily by finite $\varepsilon_{0}$.
@@
Then we will present the result of running with all sources set to 1 in section 8. We shall again display the quantities of the analytical approximation and their corresponding numerical results, as well as the deviations from the analytical approximation.
@@
Then in the next stage in section 9 we will analyze the consequences of running with finite $T_{0}$ (the value of $T=\frac{u+v}{2}$ on the initial surface) and finite $du$ – by running with several different values of these parameters and comparing the results.
@@
Afterwards in section 10 we will display that the detail of sources doesn’t matter, only the values of $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$. We will do it by running with 3 different combinations of the parameters $Z_{R0}$$Z_{Rr}$$Z_{S0}$$Z_{Sr}$ chosen such that they all lead to the same $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$. We display the metric functions $R$ and $S$ in these configurations, showing that they are the same in those 3 configurations.
@@
Lastly we display the results of running with three other different combinations of source parameters and show the validity of the analytical approximation in all those runs in section 11
Note that because we are trying to validate the analytical approximation, the numerical simulations do not incorporate the analytical approximation in their assumptions.
7 Numerical results - run with no sources
@@
We start by presenting the results of running with zero source terms, $Z_{R0}=$$Z_{Rr}$$Z_{S0}=$$Z_{Sr}=$0=0, as well as $\delta Q=0$, for $\frac{Q}{M}=0.8$. In this case, $T_{uu}=0$ throughout the whole spacetime, as already mentioned in subsection 2.8 above. We will first look at the entire domain of integration (including all three regions). Then we shall have a more focused look: first on region 2, and then on region 3. In region 2 we will compare the results with the local RN metric. In region 3 we will compare the metric functions to the analytical approximation from section 4. We will then analyze the deviation from the analytical approximation.
7.1 Overview
Figure 7.1: R
@@ 7.1 Overview
As one can observe in Figure 7.1a, showing $R(v)$, $R$ starts at a value of $R_{+}\equiv r_{+}^{2}=2.56$ in region 1, drifts down because of the negative $M_{,v}$ (still in region 1); then it “falls down” during the active region (region 2) – as it does in RN – and ends near $R_{-}\equiv r_{-}^{2}=0.16$ near IH, at the beginning of region 3. Note that this apparently rapid decrease in $R$ is actually smooth and regular. This decrease in $R$ from $r_{+}$ to $r_{-}$ happens in a timescale of order a few times $M$ (just like in RN). It only appears sharp because of the horizontal scale which spans a very large range of $v$. Then $R(v)$ drifts slightly further downwards with a small negative $R_{,v}$ according to Eq. (4.40).
The second panel 7.1b, displaying $R(u)$, shows a slightly different behavior. Because $T_{uu}=0$ (unlike $T_{vv}$ which is not zero), $R$ stays almost constant in region 1, $R\approx r_{+}^{2}\left(M(v=900),Q(v=900)\right)=0.551535$, and also in region 3 where $R\approx r_{-}^{2}\left(M(v=900),Q(v=900)\right)=0.034471$. It only varies significantly in region 2, similar to the RN case. In this special case where $J_{u}=0$ and $T_{uu}=0$ the geometry is the charged Vaidya solution, therefore along an incoming null ray with $v=const$ all quantities behave exactly like in the corresponding RN.
Figure 7.2: S
@@
The following figure 7.2 shows the behavior of $S$. We first focus on $S(v)$ in panel 7.2a. In region 1, $S_{,v}=\kappa_{+}(v)$, and hence $S$ increases towards region 2 (but this is hard to notice in the figure because of the vertical scale, which mostly shows region 3. The behavior in regions 1 and 2 is better seen in Figure 7.8a below). In region 2, $S$ peaks and then goes down as $S_{,v}=-\kappa_{-}(v)$ on approaching the near-IH region. In region 3, $S$ keeps drifting down over the long stretch from $v=100$ until $v=900$ for the same reason. $S$ eventually reaches very negative values of -4000 and even below. Therefore, $e^{S}\approx 0$ to a very good approximation.
Panel 7.2b is very similar to panel 7.2a, and shows $S$ as a function of $u$, along the other ray $v=900$.
Figure 7.3: $\widetilde{T}_{uu}$ on last v,u rays
@@
Next, we show in Figure 7.3 that $\widetilde{T}_{uu}=0$ in the entire domain of integration. The small deviation from zero (primarily seen in region 2) is due to numerical truncation error (we verified this by comparing different $du$ values).
Figure 7.4: $\widetilde{T}_{vv}$ on last u ray
@@
We continue to Figure 7.4 which shows both $\widetilde{T}_{vv}$ and the quantity
$2M_{,v}-2\frac{1}{r}QQ_{,v}$ (7.1)
which is the expected value of $\widetilde{T}_{vv}$ according to Eq. (2.107). The two curves overlap, demonstrating that $\widetilde{T}_{vv}$ matches its expected value. In region 2, $\widetilde{T}_{vv}(v)$ goes up from -0.0004 to +0.0004 (due to the rapid change in $r$ in the above expression $2\frac{1}{r}QQ_{,v}$). It then continues to drift in region 3 (though more slowly) as all factors $M_{,v}$ , $r$,$Q$, $Q_{,v}$ appearing in 7.1 slowly drift there with increasing $v$.
Figure 7.5: $\widetilde{T}_{vv}(u)$ on last v ray
Next we see in Figure 7.5 that $\widetilde{T}_{vv}(u)$ is constant in regions 1 and 3, and it only varies in region 2. This is because $M_{,v}$ and $Q_{,v}$ and $Q$ depend only on $v$, hence these quantities are constant along the last $v$ ray. Therefore the only change in the expression 7.1 comes from changing $r$. But $r$ doesn’t change in regions 1 and 3, it only changes in region 2.
Figure 7.6: $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ on initial hypersurface
@@
(b) $\widetilde{T}_{vv}-\left(2M_{,v}-2\frac{1}{r}QQ_{,v}\right)$ on the initial hypersurface
@@
In Figure 7.6 we elaborate on the fluxes $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ on the initial hypersurface $T=T_{0}$ (the red line in Figure 6.1). Panel 7.6a compares $\widetilde{T}_{vv}$ along with its theoretically expected value $2M_{,v}-2\frac{1}{r}QQ_{,v}$. $\widetilde{T}_{vv}$ is computed according to 2.42. We display the small difference between these two quantities in Figure 7.6b, which shows the error reaching at most $2\cdot 10^{-5}$. We see that the computed value matches the expected value well. Figure 7.6c, shows that up to numerical errors ($2\cdot 10^{-10}$, mainly roundoff), $T_{uu}=0$ on the initial hypersurface line, as it should.
7.2 Region 2: comparison with the local RN approximation